# Fast nonparametric feature selection with error control using integrated path stability selection

### v1

@@ 2.  Methods
is the impurity difference between $v$ and its children, $v_{L}$ and $v_{R}$. Large positive values of $\Delta\varphi(v)$ indicate that the feature $j_{v}$ used to split node $v$ successfully partitions the data in a manner that is consistent with the objective of the tree. Importance scores naturally lend themselves to thresholding: Given a $\Phi$ and a threshold $\lambda$, the selected set of features is $\hat{S}_{\lambda}=\{j:\Phi(j)\geq\lambda\}$.
Algorithm S1 outlines IPSS for thresholding. The main steps are as follows: First, features are preselected according to the procedure described in Section S1.4. This is a common preliminary step in many feature selection algorithms. Next, importance scores are evaluated for the preselected features using random, disjoint halves of the data. This process is repeated $B$ times, yielding $2B$ importance scores for each feature. We then construct a grid of $\lambda$ values. The largest, $\lambda_{\mathrm{max}}$, is the maximum importance score across all features and all $2B$ sets of scores (hence, $\hat{S}_{\lambda_{\mathrm{max}}}=\varnothing$). Next, starting from $\lambda_{\mathrm{min}}=\lambda_{\mathrm{max}}$, we decrease $\lambda_{\mathrm{min}}$ one step at a time, usually on a log scale, iteratively updating the integral $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ at each step in the form of a Riemann sum approximation until $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ surpasses the cutoff $C=0.05$. Once $C$ is surpassed, the $\mathtt{while}$ loop stops and the feature-specific selection probabilities and integrals are evaluated. The algorithm outputs the efp scores for each feature. Given a target E(FP), $t$, the set of selected features is $\hat{S}(t)=\{j:\mathtt{efp}_{Z_{1:n}}(j)\leq t\}$.
2.4.  IPSSGB and IPSSRF.
We consider two instances of the general method: IPSSGB uses gradient boosting for the baseline algorithm $\hat{S}_{\lambda}$, and IPSSRF uses random forests for $\hat{S}_{\lambda}$. All other IPSS parameters are identical, except we use $B=50$ for IPSSRF and $B=100$ for IPSSGB. The latter led to slightly more stable results with hardly any increase in runtimes (no such improvement was observed for IPSSRF).
In all that follows, random forests are implemented with scikit-learn (Pedregosa et al., 2011) and gradient boosting with XGBoost (Chen and Guestrin, 2016). For random forests, all parameters are set to their default settings except for the proportion of features considered when splitting each node (called $\mathtt{max\_features}$ in scikit-learn and often $\mathtt{mtry}$ elsewhere), which we change from $1$ to $1/3$. This is a standard choice (Biau and Scornet, 2016) and led to an approximately $3$-fold speedup in IPSSRF without any discernible effect on the results. For gradient boosting, we also change the proportion of features considered at each split from $1$ to $1/3$. Aside from this, the only non-default parameter is the maximum depth of each tree, which we change from $6$ to $1$, making each tree a stump. This significantly improved the performance of IPSSGB, both in terms of speed and feature selection. Finally, while many different importance functions exist for gradient boosting and random forests, we use mean decrease impurity, Equation 2.4, with the squared error loss for regression and the Gini index for classification, as defined in Section 2.3. Experimentation with other importance functions showed similar results, though a more detailed study of IPSS with different importance functions is a potentially direction for future work.
2.4.1.  Further background.
The seminal work on gradient boosting is Friedman (2001). For random forests, we refer to the original work of Breiman (2001), and to Biau and Scornet (2016) for an excellent review with a detailed description of importance scores. Gradient boosting, random forests, and their importance scores are also detailed in Hastie et al. (2009). Finally, our formulation of tree-based scores partly follows Louppe et al. (2013), who provide a clear and concise formulation of these quantities.
3.  Other methods
We describe several feature selection methods that are compared to IPSSGB and IPSSRF in Sections 4 and 5. We selected these methods due to their popularity and superior performance relative to other methods in several comparison studies (Degenhardt et al., 2019; Sanchez-Pinto et al., 2018; Speiser et al., 2019).
3.1.  Methods with error control.
The closest method to IPSSGB in terms of its underlying approach is that of Hofner et al. (2015), referred to here as SSBoost. Unlike IPSSGB—which uses importance scores from gradient boosting—SSBoost applies stability selection to choose the number of features used per boosting run. Furthermore, IPSSGB uses IPSS to construct efp scores, whereas SSBoost uses a version of stability selection introduced by Shah and Samworth (2013). This is perhaps the most significant difference since the efp scores for IPSS have much tighter bounds than those for other forms of stability selection (Melikechi and Miller, 2024). From a practical standpoint, this causes other versions of stability selection to identify fewer important features than IPSS for the same target E(FP).
IPSS has previously been used by Melikechi and Miller (2024) with lasso for regression and $\ell^{1}$-regularized logistic regression for classification (Tibshirani, 1996; Lee et al., 2006). These methods, abbreviated as IPSSL, are part of a more general class of $\ell^{1}$-regularized estimation techniques to which stability selection is commonly applied (Meinshausen and Bühlmann, 2010), namely,
$\hat{\beta}$ $=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\sum_{i=1}^{n}\mathcal{L}(Y_{i},\beta^{\mathtt{T}}X_{i})+\lambda\sum_{j=1}^{p}\lvert\beta_{j}\rvert,$
where $\mathcal{L}$ is a log-likelihood function and $\lambda>0$ controls the strength of the $\ell^{1}$ penalty. By contrast, since decision trees are nonparametric, neither IPSSGB nor IPSSRF assume an underlying model. Furthermore, unlike linear regression and logistic regression, decision trees can effectively capture nonlinear relationships in data. The nonlinear and nonparametric nature of decision trees motivated our extension of IPSS to gradient boosting and random forests. In Section 4, we find that both methods significantly outperform IPSSL in the presence of nonlinearity.
Another method to which we compare is knockoff boosted trees (KOBT) (Jiang et al., 2021). KOBT applies the model-X knockoff framework of Candes et al. (2018) to boosted trees to obtain a method with theoretical FDR control. We use the R package KOBT, preselecting $\mathtt{bound}=200$ features and generating $\mathtt{num}=100$ knockoff matrices. The knockoff type is shrunken Gaussian. These choices were made to balance performance and runtime. Computation times were much longer for larger values of $\mathtt{bound}$ or $\mathtt{num}$, or when using a different version of KOBT that employs principal components to avoid the Gaussian assumption. Moreover, we found that these computationally more expensive choices did not noticeably improve performance or FDR control.
We also tested the method of Coleman et al. (2022), which achieves theoretical error control by using random forests for hypothesis testing. However, one test run with default parameters on simulated data with 500 samples, 500 features, and 20 true features took over 51 minutes, returning 15 true positives and 43 false positives. By contrast, IPSSGB with a target E(FP) of 2 took 11 seconds and returned 10 true positives and 0 false positives on the same data. This method is omitted from our study because its performance does not appear to justify its excessive runtime.
3.2.  Methods without error control.
We also compare to several popular methods without theoretical error control: Boruta (Kursa et al., 2010), random forests (RF) (Breiman, 2001), Vita (Janitza et al., 2018), VSURF (Genuer et al., 2010), and XGBoost (Chen and Guestrin, 2016). Boruta, Vita, and VSURF use random forests. In terms of performance, Speiser et al. (2019) found Boruta and VSURF were among the best out of $13$ random forest-based methods, and Degenhardt et al. (2019) found Boruta and Vita outperformed $5$ other methods on several metrics.
Method Package Error control Base method Non-default settings
IPSSGB ipss (Python) Boosting
IPSSRF ipss (Python) Random forest
IPSSL ipss (Python) Lasso
SSBoost XGBoost (Python) Boosting $\tau=0.75$
with stabs (R)
KOBT KOBT (R) Boosting $\texttt{num}=100$, $\texttt{bound}=200$,
[…]
Boruta Boruta (R) Random forest
RF scikit-learn (Python) Random forest
Vita vita (R) Random forest $p$-value threshold $=0$
VSURF VSURF (R) Random forest VSURF_pred
XGBoost XGBoost (Python) Boosting
Table 1. Feature selection methods. Packages, when available, are listed with the language used to implement them in parentheses. Non-default settings are elaborated upon in Sections 3.1 and 3.2. For methods with no non-default settings, we use the default settings in their respective packages.
Table 1 summarizes the methods. Details regarding the implementation and choice of settings are as follows. Hofner et al. (2015) provide code for SSBoost that combines the R packages mboost (Hofner et al., 2014) and stabs (Hofner and Hothorn, 2017) in the form of a worked example, but the mboost implementation of boosting was prohibitively slow in the dimensions we consider. By adapting their code to use XGBoost in place of mboost for boosting and by preselecting features exactly as we do for IPSSGB (Section S1.4), we were able to reduce SSBoost runtimes considerably with no apparent change in results. The XGBoost parameters used for SSBoost are the same as those used for IPSSGB; specifically, all XGBoost parameters are set to their default values except the proportion of features considered at each split (changed from 1 to 1/3) and the depth of each tree (changed from 6 to 1). For the stability selection part of SSBoost we use the default parameters in stabs, and the selection threshold is set to $\tau=0.75$, which is the middle of the interval $(0.6,0.9)$ recommended by Meinshausen and Bühlmann (2010). For Vita, we set the $p$-value threshold to 00, and for VSURF we use the function VSURF_pred rather than VSRUF_interp to select the final set of features (Genuer et al., 2010); both choices correspond to as few features as possible being selected under the constraints of the respective algorithms, a favorable quality due to the sparsity in our forthcoming studies. For XGBoost and RF, we run the respective algorithms on the full dataset with default parameters, then select the top ranked features by their importance scores.
4.  Simulation study
In this section, we conduct a simulation study to evaluate the performance of the methods in Table 1 when the true set of important features is known. To make the simulations more realistic, we use features from real data rather than generating them from known distributions. Specifically, given $n$ and $p$, in each trial we randomly select $n$ samples and $p$ features from RNA sequencing (RNA-seq) data from 569 ovarian cancer patients (samples) and 6426 genes (features) (Vasaikar et al., 2018). This publicly available dataset, part of The Cancer Genome Atlas (Weinstein et al., 2013), was chosen because it is high dimensional and the features follow a variety of distributions. For instance, Figure 2 shows the standardized empirical distributions of five randomly selected genes, which, from left to right, are relatively flat, skewed left, approximately Gaussian, skewed right, and contain outliers. Furthermore, the genes exhibit complex correlation structures, with maximum and average absolute pairwise correlations of approximately $0.95$ and $0.17$ after standardization, respectively.
[…]
Figure 1. Simulating the response. Random rows and columns are selected from the full RNA-seq dataset to create an $n$-by-$p$ design matrix $X$, whose columns are then split into important features, $X_{S}$, and unimportant features, $X_{S^{c}}$. The columns of $X_{S}$ are further partitioned into $G$ groups; the above figure shows the case $G=3$. The features in each group are summed to obtain $\xi_{g}$, and a different realization $f_{\theta_{g}}$ of $f_{\theta}$ is applied to $\xi_{g}$ for each $g$. The response is the sum of the resulting group-specific signals, plus noise.
Figure 2. Distributions of five genes from the RNA-seq dataset.
4.1.  Simulation details.
Algorithm S2 describes the simulation procedure, which is also illustrated in Figure 1. In all steps, “randomly select” means select a parameter uniformly at random from its domain. The general outline is as follows: First, randomly select an $n$-by-$p$ submatrix $X$ of the full RNA-seq dataset and standardize its columns to have mean 0 and variance 1. Next, randomly select a true subset of $p_{\mathrm{true}}$ important features $S$, and partition it into $G$ groups, $S_{1},\dots,S_{G}$. A different realization of a randomized nonlinear function $f_{\theta_{g}}$, defined in Equation S2.1, is applied to the standardized sum $\xi_{g}=(\hat{\xi}_{g}-\bar{\xi}_{g})/\bar{\sigma}_{g}$ of the features in each group $S_{g}$, where $\bar{\xi}_{g}$ and $\bar{\sigma}_{g}$ are the empirical mean and standard deviation of $\hat{\xi}_{g}=\sum_{j\in S_{g}}X_{j}$. The resulting values are summed over all groups to generate a signal $\eta=\sum_{g=1}^{G}f_{\theta_{g}}(\xi_{g})$, and noise is added to this signal to generate a response $y$. This scheme produces data with highly complex interactions between features and the response, going well beyond the additive setting $y=\sum_{j\in S}f_{j}(X_{j})+\epsilon$.
@@ 4.  Simulation study
Regression and classification data are simulated with $p=500$, $2000$, and $5000$ features for a total of six experiments. Each experiment consists of $100$ trials, outlined in Algorithm S2. Each trial uses $n=500$ samples and $\lvert S\rvert=p_{\mathrm{true}}$ important features. For regression, the response is $y=\sum_{g=1}^{G}f_{\theta}(\xi_{g})+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ and the variance $\sigma^{2}$ is selected according to a specified signal-to-noise ratio (SNR). For classification, we draw $y_{i}\sim\mathrm{Bernoulli}(\pi_{i})$ where $\pi_{i}=1/(1+\exp(-u\eta_{i}))$. A new $p_{\mathrm{true}}$ and SNR (or $u$), as well as new sets of samples, features, and important features, are drawn prior to each trial, with $p_{\mathrm{true}}$, SNR, and $u$ chosen according to Table S1. The many sources of randomness in our setup are introduced so that the simulations cover a wide range of settings.
4.2.  Simulation results.
Figures 3, LABEL:, 4, LABEL:, and 5 show the results. The empirical false discovery and true positive rates are $\mathrm{FDR}=\mathrm{FP}/(\mathrm{TP}+\mathrm{FP})$ and $\mathrm{TPR}=\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})$, where FP, TP, and FN are the number of false positives, true positives, and false negatives, respectively. The upper left (high TP or TPR and low FP or FDR) corresponds to better performance in Figures 3 and 4. The ability of the different methods to control E(FP) (Figure 3) and FDR (Figure 4) at their target values is shown when applicable. Specifically, SSBoost and KOBT provide theoretical control of E(FP) and FDR, respectively, while the IPSS methods can be specified to control either E(FP) or FDR (Section 2.2). In Figure 3, the dashed vertical line signifies a target E(FP) of 3, with symbols showing the average FP and TP for the relevant methods at this target value. In particular, the for a method with perfect E(FP) control should lie on the dashed vertical line. Similarly, the and symbols in Figure 4 correspond to target FDRs of 0.1 and 0.3, respectively.
IPSSGB has the best performance overall. Its curves are the farthest to the upper left in every plot in Figures 3 and 4, and it has the best control of both E(FP) and FDR. In Figure 3, for example, the for IPSSGB is the closest to the vertical dashed line in all experiments, and similarly in Figure 4. Furthermore, IPSSGB identifies the most true positives, both in terms of average TP and average TPR. Notably, at a target E(FP) of 3 and a target FDR of 0.3, IPSSGB has the largest average TP and highest TPR in nearly every setting, while having significantly fewer false positives than most other methods. Figure 5 also shows that IPSSGB is one of the fastest methods, running in under 20 seconds in all settings.

### v2

@@ 2.  Methods
is the impurity difference between $v$ and its children, $v_{L}$ and $v_{R}$. Large positive values of $\Delta\varphi(v)$ indicate that the feature $j_{v}$ used to split node $v$ successfully partitions the data in a manner that is consistent with the objective of the tree.
For regression, we use the squared error loss impurity function, $\varphi(v)=\sum_{i\in v}(Y_{i}-\overline{Y}_{v})^{2}/\lvert v\rvert$, where $v\subseteq\{1,\dots,n\}$ is identified with the subset of samples in node $v$, and $\overline{Y}_{v}=\sum_{i\in v}Y_{i}/\lvert v\rvert$ is the empirical mean of the responses in $v$. For binary classification, we use the Gini index, $\varphi(v)=2p_{0}(v)p_{1}(v)$ where, for $a\in\{0,1\}$, $p_{a}(v)=\sum_{i\in v}\operatorname{\mathds{1}}(Y_{i}=a)/\lvert v\rvert$ is the proportion of responses in $v$ that equal $a$. These were selected because they are canonical choices and because we found little difference in results between these and other standard impurity functions.
The importance function defined in Equation 2.5 can be computed for any tree ensemble method. We focus on gradient boosting (Friedman, 2001) and random forests (Breiman, 2001). As noted above, these are abbreviated IPSSGB and IPSSRF when combined with IPSS. Additional implementation details about IPSSGB and IPSSRF are in Sections S2.2.1 and S2.2.2, respectively.
2.3.3  Other importance functions.
The importance function in Equation 2.5 is called mean decrease impurity (MDI) because it measures the average decrease in impurity attributed to each feature over all trees. Another common importance function is mean decrease accuracy (MDA), also known as permutation importance (Louppe et al., 2013).
MDA is more general than MDI in that it can be applied to any predictive model, not just tree ensembles. Several variants of MDA exist, but the basic procedure is: (i) train the model, (ii) compute the prediction error $e$ on test data, and (iii) for each feature $j$, randomly permute the values of feature $j$ in the test data (keeping all other features unchanged), and compute the prediction error $e_{j}$ of the trained model on the permuted test data. The underlying idea, captured by the importance score $\Phi(j)=e_{j}-e$, is that permuting unimportant features should have little effect on the prediction error, while permuting important features should degrade predictive performance.
We experimented with IPSS applied to both MDI and MDA from gradient boosting and random forests. In both cases, the FDR was controlled at target levels, but IPSS with MDI consistently identified more true positives. One likely reason for this is that MDA tends to spread importances more uniformly across features than MDI, making it more difficult to distinguish between important and unimportant features (Hastie et al., 2009). Furthermore, MDI, whose importance scores are obtained during the training process itself, is more computationally efficient than MDA, which requires the additional step of evaluating the trained model on the permuted data for every feature.
The recent success of deep learning, especially on text and image data, has generated much interest in feature importance scores—for example, SHAP values (Lundberg and Lee, 2017)—that are derived from these models. Applying IPSS to such scores is a potentially interesting line of future work. However, numerous studies have shown that tree ensemble methods like XGBoost (Chen and Guestrin, 2016) are faster, easier to train, and consistently outperform deep learning methods on tabular data, especially when there are fewer than 10,000 samples (Shwartz-Ziv and Armon, 2022; Grinsztajn et al., 2022; Fayaz et al., 2022). Since our focus in this work is on tabular medical data, where the number of patients rarely exceeds several hundred, we primarily consider importance scores from tree ensemble methods given their many advantages over deep learning in this setting.
2.3.4  Computation.
The IPSS subsampling procedure requires $2B$ feature importance evaluations, one evaluation on each half of the data in all $B$ iterations. This is fast for MDI scores; when combined with preselection (Section S2.1), IPSSGB and IPSSRF run in under 20 seconds on a standard laptop when there are $n=500$ samples and $p=5000$ features (Tables S3 and S4). Since this is already sufficiently fast for our purposes, we do not implement more efficient alternatives in this work.
We note, however, that IPSS is embarrassingly parallel: All iterations of the subsampling procedure can be evaluated separately and hence run simultaneously, potentially accelerating IPSS by a factor of $2B$ (typically 100 or 200). This is especially beneficial when importance scores are expensive to compute, as is often the case when they come from deep learning methods. By contrast, many feature selection methods are iterative and thus not parallelizable. For example, recursive feature elimination (Li et al., 2022) typically removes features one at a time based on predictive performance, requiring the model to be retrained after each removal. When $p$ is large, this can require thousands of model fits, far more than the $2B$ evaluations typically required by IPSS.
3.  Simulation studies
In this section, we conduct two simulation studies to evaluate the performance of 14 feature selection methods when the true set of important features is known. Features in the first study are drawn from a multivariate Gaussian, and the response is generated from a nonlinear additive model. To make the simulations more realistic, in the second study we use features from real RNA sequencing (RNA-seq) data rather than generating them from known distributions, and the response is a highly randomized, nonlinear function of the important features.
3.1  Other methods.
We compare IPSSGB and IPSSRF to 12 feature selection methods: IPSS with $\ell^{1}$-regularization (IPSSL1, Melikechi and Miller (2024)); boosting with stability selection (SSBoost, Hofner et al. (2015)); five versions of model-X knockoffs (Candes et al., 2018), namely knockoffs with generalized linear models (KOGLM), knockoffs with lasso (KOL1), knockoffs with random forests (KORF), knockoffs with boosted trees (KOBT, Jiang et al. (2021)), and knockoffs with deep neural networks (DeepPINK, Lu et al. (2018)); random forest hypothesis testing (RFHT, Coleman et al. (2022)); Boruta (Kursa et al., 2010); recursive feature elimination with gradient boosting (RFEGB); Vita (Janitza et al., 2018); and VSURF (Genuer et al., 2010). We provide a brief overview of these methods below. Additional details, including parameter settings and the software packages used to implement each method, are in Section S2.
Four methods—Boruta, RFEGB, Vita, and VSURF—do not have theoretical false discovery control. We chose these because Speiser et al. (2019) found that Boruta and VSURF were among the best out of $13$ random forest-based feature selection methods, and Degenhardt et al. (2019) found that Boruta and Vita outperformed $5$ other methods in extensive comparison studies. We tested RFEGB so as to include a gradient boosting-based feature selection method without false discovery control.
IPSSL1 and SSBoost provide theoretical E(FP) control. IPSSL1, a parametric version of IPSS, assumes a (generalized) linear relationship between the features and the response (Equation 2.4). SSBoost combines gradient boosting and stability selection. We implement it assuming the $r$-concavity conditions of Shah and Samworth (2013), which are required to obtain the tightest upper bound on E(FP) of any version of stability selection other than IPSS (note that IPSS does not require $r$-concavity assumptions) (Melikechi and Miller, 2024).
Model-X knockoffs is a general framework for feature selection with theoretical FDR control that has attracted much attention recently, in part due to its flexibility. Like IPSS, model-X knockoffs works in high-dimensions ($p>n$), does not use $p$-values (which are challenging to compute in general), is compatible with arbitrary feature importance scores, and makes no assumptions about the relationship between the response and the features. Unlike IPSS, model-X knockoffs assumes that the joint distribution of the features is known (Candes et al., 2018).
3.2  Gaussian simulation design.
We perform two regression experiments and two classification experiments. The two experiments in both settings correspond to $n=250$ and $500$. All experiments consist of $100$ trials. In each trial, $n$ independent samples of $X\sim\mathcal{N}(0,\Sigma)$ are drawn from a $p=500$-dimensional, mean-zero multivariate Gaussian with a Toeplitz covariance matrix, $\Sigma$. The correlation parameter of the Toeplitz matrix is set to $\rho=0.5$; that is, $\Sigma_{jk}=0.5^{\lvert j-k\rvert}$.
@@ 3.  Simulation studies
The number of important features $\lvert S\rvert$ is drawn uniformly at random from $\{5,\dots,15\}$ prior to each trial, and a new set of important features $S$ of size $\lvert S\rvert$ is randomly selected. In each experiment, the signal is $f(X)=\sum_{j\in S}\exp(-X_{j}^{2})$. For regression, the response is $Y=f(X)+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ and $\sigma^{2}$ is selected according to a specified signal-to-noise ratio (SNR) that is drawn uniformly at random from the interval $[0.5,2]$. For classification, we draw $Y\sim\mathrm{Bernoulli}(\pi)$ where $\pi=1/(1+\exp(-uf(X)))$ and the signal strength $u$ is drawn uniformly at random from the interval $[1,3]$. These sources of randomness are introduced so that the study covers a wide range of settings.
Figure 1. Gaussian simulation results ($n=500$). First and second columns show the regression and classification results, respectively. The horizontal axis in each plot shows the target FDR. The three methods without false discovery control (gray horizontal lines) do not vary with the target FDR. The black dashed line in the FDR plots represents perfect FDR calibration, FDR = target FDR. Non-gray solid lines correspond to IPSS or stability selection-based methods. Non-gray dashed lines show methods based on model-X knockoffs.
3.3  RNA-seq simulation design.
We perform three regression experiments and three classification experiments. The three experiments in both settings correspond to $p=500$, $2000$, and $5000$. All experiments consist of $100$ trials. In each trial, $n=500$ patients and $p$ genes are randomly selected from RNA-seq measurements of 6426 genes from $569$ ovarian cancer patients (Vasaikar et al., 2018). This publicly available dataset, part of The Cancer Genome Atlas (Weinstein et al., 2013), was chosen because it is high dimensional and the features follow a variety of empirical distributions (Figure S5). Furthermore, the genes exhibit complex correlation structures, with maximum and average absolute pairwise correlations of approximately $0.95$ and $0.17$ after standardization, respectively.
Algorithm S2 describes the simulation procedure for each individual trial, which is also illustrated in Figure S3. In all steps, “randomly select” means select a parameter uniformly at random from its domain. The general outline is as follows: First, randomly select an $n$-by-$p$ submatrix $X$ of the full RNA-seq dataset and standardize its columns to have mean 0 and variance 1. Next, the number of important features $\lvert S\rvert$ is drawn uniformly at random from $\{10,\dots,30\}$ and a randomly selected subset of $\lvert S\rvert$ important features $S$ is partitioned into $G$ groups, $S_{1},\dots,S_{G}$. A different realization of a randomized nonlinear function $f_{\theta_{g}}$, defined in Equation S3.1, is applied to the standardized sum $\xi_{g}=(\hat{\xi}_{g}-\bar{\xi}_{g})/\bar{\sigma}_{g}$ of the features in each group $S_{g}$, where $\bar{\xi}_{g}$ and $\bar{\sigma}_{g}$ are the empirical mean and standard deviation of $\hat{\xi}_{g}=\sum_{j\in S_{g}}X_{j}$. The resulting values are summed over all groups to generate a signal $\eta=\sum_{g=1}^{G}f_{\theta_{g}}(\xi_{g})$, and noise is added to this signal to generate a response $Y$. This scheme produces data with highly complex interactions between features and the response, going well beyond the additive setting $Y=\sum_{j\in S}f_{j}(X_{j})+\epsilon$.
For regression, the response is $Y=\eta+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ and the variance $\sigma^{2}$ is selected according to a specified signal-to-noise ratio (SNR) that is drawn uniformly at random from the interval $[0.5,2]$. For classification, we draw $Y\sim\mathrm{Bernoulli}(\pi)$ where $\pi=1/(1+\exp(-u\eta))$ and the signal strength $u$ is drawn uniformly at random from $[1,3]$. As in the Gaussian simulation experiments, these many sources of randomness are introduced to cover a wide range of settings.
3.4  Simulation results.
Figures 1 and S1 show the results of the Gaussian simulations when $n=500$ and $250$, respectively, and Figure 2 and S2 show the results of the RNA-seq simulations for regression and classification. Runtimes for each method in each experiment are provided in Tables S3 and S4. The FDR in each plot is the average of $\mathrm{FP}/(\mathrm{TP}+\mathrm{FP})$ over all 100 trials, and the true positive rate (TPR) is the average of $\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})$, where FP, TP, and FN are the number of false positives, true positives, and false negatives, respectively.
Both FDR and TPR are shown as functions of the target FDR. The black dashed line in each FDR plot represents perfect FDR calibration, $\mathrm{FDR}=\mathrm{target\ FDR}$. A method’s FDR is well-controlled if its FDR lies on or below this line. The FDR and TPR for methods without false discovery control are shown as horizontal lines because they do not admit false discovery control parameters and therefore do not vary with the target FDR.
IPSSGB has the best performance overall. Its FDR is always well-controlled and it consistently has a much higher TPR than the other methods with false discovery control. Notably, IPSSGB identifies significantly more true positives than other methods with false discovery control at lower target FDRs. For example, in the regression results in Figure 1, IPSSGB identifies 70% of important features when the target FDR is 0.1, while IPSSRF identifies 40% and the remaining methods identify close to none. IPSSGB’s TPR even surpasses the TPR of methods without error control in almost all experiments despite having far fewer false positives. With an average runtime of less than 15 seconds in all experiments, IPSSGB is also one of the fastest methods.